OpenStudy (anonymous):

Wallace took out a $5,000 loan for six years. He is being charged 4 percent interest, compounded annually. Calculate the total amount he will pay.

OpenStudy (anonymous):

A=P(1+r/n)^nt A=5000(1+.04/1)^1*6 A=5000(1.04)^6 A=5000(1.265319018) A=6,326.60 (rounded)